OpenStudy (anonymous):

What is the kinetic energy of a 1.0-kg billiard ball that moves at 5.0 m/s? A. 5.0 J B. 6.5 J C. 10.0 J D. 12.5 J

563blackghost (563blackghost):

EK= m x(times) V2(squared)/2. So....1x5x5/2=? I can tell you already that A and B are out

563blackghost (563blackghost):

sorry not squared its square

563blackghost (563blackghost):

So 1x5x5/2=? 5x5 /2= 25/2=

563blackghost (563blackghost):

hope this helps please leave a medal! xD